PROBLEM 8-24 Cash Budget with Supporting Schedules [LO2, LO4, LO8) Garden Sales, Inc., sells garden supplies. Management is planning its cash needs for the second quarter. The company usually has to borrow money during this quarter to support peak sales of lawn care equip- ment, which occur during May. The following information has been assembled to assist in preparing a cash budget for the quarter: a. Budgeted monthly absorption costing income statements for April-July are: April $800,000 420,000 180,000 Cost of goods sold Gross margin Seling and administrative expenses: $900,000 630,000 $500,000 350,000 150,000 $400,000 280,000 120,000 270000 Seling expense 79,000 45,000 120,000 52,000 62.000 41,000 51,000 38,000 Administrative expense Total seling and administrative. Net operating income Includes $20,000 of depreciation each month 124,000 172,000 103,000 89,000 $56.000-8,000 $47,000 $31,000 b. Sales are 20% for cash and 80% on account Sales on account are collected over a three-month period with 10% collected in the month of sale; 70% collected in the first month following the month of sale; and the remaining 20% collected in the second month following the month of sale. Februarys sales totaled $200,000, and Marchs sales totaled $300,000 Inventory purchases are paid for within 15 days. Therefore, 50% of a months inventory purchases are paid for in the month of purchase. The remaining 50% is paid in the following month. Accounts payable at March 31 for inventory purchases during March total S126,000 Each months ending inventory must equal 20% of the cost of the merchandise to be sold in the following month. The merchandise inventory at March 31 is $84,000. Dividends of $49,000 will be declared and paid in April Land costing $16,000 will be purchased for cash in May The cash balance at March 31 is $52,000; the company must maintain a cash balance of at least $40,000 at the end of each month The company has an agreement with a local bank that allows the company to borrow in increments of $1,000 at the beginning of each month, up to a total loan balance of $200,000. The interest rate on these loans is 1% per month and for simplicity we will assume that interest is not compounded. The company would, as far as it is able, repay the loan plus accumulated interest at the end of the quarter. e. d. e. f. g. h. i. Required I. 2. Prepare a schedule of expected cash collections for April, May, and June, and for the quarter in total. Prepare the following for merchandise inventory: a. A merchandise purchases budget for April, May, and June A schedule of expected cash disbursements for merchandise purchases for April, May, and June, and for the quarter in total. h 3. Prepare a cash budget for April, May, and June as well as in total for the quarter
Answer #1

1.

Schedule of Expected Cash Collection
April ($) May ($) June ($) Quarter
Cash Sale in each month
(20% of total sales)
         120,000          180,000          100,000              400,000
Collection from on Account Sales (See Working Notes)          248,000          456,000          640,000          1,344,000
   368,000.00    636,000.00    740,000.00    1,744,000.00
Working Note -1
Collection in the month of April
10% of April's on Account Sale =(600000*80%)*10% 48000
70% of March's on Account Sale =(300000*80%)*70% 168000
20% of Feb's on account Sale =(200000*80%)*20% 32000

248000

Working Note -2
Collection in the month of May
10% of May's on Account Sale =(900000*80%)*10% 72000
70% of April's on Account Sale =(600000*80%)*70% 336000
20% of March's on account Sale =(300000*80%)*20% 48000
456000
Working Note -2
Collection in the month of June
10% of June's on Account Sale =(500000*80%)*10% 40000
70% of May's on Account Sale =(900000*80%)*70% 504000
20% of April's on account Sale =(600000*80%)*20% 96000
640000